# Client Selection for Federated Policy Optimization with Environment Heterogeneity

### v5

@@ 3.4 Connection to Centralized Learning
When the environments are homogeneous, the learning process of FAPI will be exactly the same as the learning from N copies of the same environment (which is identical to the imaginary MDP) with federated policy evaluation. Under such circumstances, Theorem 10 degenerates to that for centralized learning (6) (Bertsekas, 2022) with the same error bound, i.e., $\frac{\bar{\epsilon}-2\gamma\bar{\delta}}{(1-\gamma)^{2}}$, as shown in the following proposition.
@@
$\limsup_{t\to\infty}\left\|\bar{V}^{\pi^{t}}-\bar{V}^{\pi^{\ast}}\right\|$ $\leq\frac{\bar{\epsilon}-2\gamma\bar{\delta}}{(1-\gamma)^{2}}.$
See Appendix F for the detailed proof.
@@ 4.2 Implementation
To compute $\Delta_{n}$, we adopt the tabular maximum likelihood model (Moerland et al., 2020; Strehl et al., 2009; Ornik and Topcu, 2021; Strehl and Littman, 2008) in the implementation of FedPOHCS. Given a set of trajectories, let $C_{n}(s,a)$ denote the number of times action $a$ was taken under state $s$, $C_{n}(s,a,s^{\prime})$ represent the number of times the MDP transited from state $s$ to $s\prime$ after taking action $a$, and $r_{n}$ denote the corresponding sequence of reward received. Generally speaking, the more trajectories we have, the more accurate the modeling would be. For Mountain Cars and Hoppers, the number of trajectories is 200. We can estimate the transition probability $P_{n}$ of the $n$-th MDP and the reward function $\mathcal{R}$ by
@@
where the time horizon $t$ is a hyperparameter. Then, we diagonalize the vector $D_{n,s^{\prime}}$ to obtain the state visitation frequency matrix $\hat{\mathbf{D}}_{\pi,\mu,P_{n}}$.
@@ 5 Experiments
In this section, we introduce two federated environments for empirical evaluation. Then, we evaluate the proposed client selection method from different aspects. All experimental results are completely reproducible and can be accessed from the GitHub repository: https://github.com/ShiehShieh/FedPOHCS.
@@ 5.1 Environments
We evaluate the effectiveness of the proposed client selection algorithm on a federated version of the mountain car continuous control problem (Moore, 1990) and the Mujoco Hopper problem (Coulom, 2002). We construct the federated environment as follows:
@@
Hoppers consists of 60 equally weighted Hopper-v3 environments developed by OpenAI Gym. The environment consists of an 11-dimensional continuous state space and a 3-dimensional continuous action space. We introduce the heterogeneity into this system by following Jin et al. (2022), i.e., the leg size is uniformly sampled from $[0.01,0.07]$, $[0.01,0.10]$ and $[0.01,0.15]$ for low-level, medium-level and high-level heterogeneity, respectively.
@@ 5.2 Experiment Settings
We use neural networks to represent policies as in Schulman et al. (2015); Vinitsky et al. (2018). Specifically, we use Multilayer Perceptrons (MLPs) with $\tanh$ non-linearity and hidden layers (64, 64). We use the SGD optimizer with a momentum of 0.9 and learning-rate decay of 0.98 and 0.9 per round for Mountain Cars and Hoppers, respectively. Hyperparameters are carefully tuned so that they are near-optimal for each algorithm. See Appendix G for more experimental details.
We compare FedPOHCS with biased and unbiased client selection methods, including FedAvg (random selection), Power-of-Choice, and GradientNorm (Chen et al., 2021; Marnissi et al., 2021; Chen et al., 2022). FedAvg randomly selects K clients. Power-of-Choice utilizes the aforementioned two-phase scheme and selects candidate clients with the highest loss (or lowest advantages/values in case of RL problem). GradientNorm selects candidate clients with the largest gradient norm. Other than the client selection scheme, all algorithms follow the same procedure described at the beginning of Section 2.1. In particular, clients perform local training with the algorithm proposed by Schulman et al. (2017), with adaptive KL penalty term. At the end of every round, the server broadcasts the global policy to all clients, order them to interact with their MDPs for several episodes (10 for Mountain Cars and 100 for Hoppers) and report the mean returns to evaluate the performance. Each experiment is averaged across three independent runs with different random seeds and parameter initializations, both of which are shared among all algorithms for a fair comparison. Confidence intervals are also reported.
Figure 1: Comparison of FedAvg, Power-of-Choice, FedPOHCS and GradientNorm on Mountain Cars with medium-level of heterogeneity. For FedAvg, the learning rate is 0.005 and the KL target is 0.003. For Power-of-Choice, GradientNorm and FedPOHCS, the learning rate is 1e-3 and the KL target is 0.003.
Figure 2: Comparison of FedAvg, Power-of-Choice, FedPOHCS, and GradientNorm on Hoppers with medium level of heterogeneity. For all algorithms, the learning rate is 0.03, the learning rate decay is 0.9 and the KL target is 0.003.
@@ 5.3 Performance and Stability
Although the original mountain car problem is simple and most modern RL algorithms can easily obtain a score over 90.0, the federated setting imposes great difficulties in solving it. Figure 1 shows the performance comparison between FedPOHCS and several baselines on Mountain Cars with medium-level heterogeneity. It can be observed that FedPOHCS has a faster convergence speed and a more stable learning process. To compute the selection metric, we discretize the state and action by rounding them off to the nearest 0.1, resulting in about 8000 states and 100 actions that are frequently visited. In each round, the first phase of FedPOHCS take 1-10 seconds and the local training take 20-30 seconds. Although the running time of the first phase may vary depending on the implementation, FedPOHCS outperforms all baselines in terms of the number of rounds and wall-clock time with our setting. In Figure 1, we have also included the communication-efficient variant, i.e., the one-phase scheme, for comparison purposes. It can be observed that the one-phase scheme may slow down and destabilize learning in the early stage of training, but the final performance is comparable to the two-phase scheme.
We can draw a similar conclusion on Hoppers with medium-level heterogeneity. As shown in Figure 2, FedPOHCS can obtain an accumulated reward of 1450 within 20 rounds of training while it takes about 80 rounds for Power-of-Choice to reach 1450, demonstrating the advantage of the proposed FedPOHCS algorithm in speeding up convergence.
@@ 6 Conclusion
This appendix contains the proof of Lemma 5 in Appendix A, the proof of Lemma 6 in Appendix B, the proof of Proposition 7 in Appendix C, the proof of Theorem 10 in Appendix D, the proof of Proposition 9 in Appendix E, the proof of Proposition 12 in Appendix F, and additional experiment setting in Appendix G.
@@ Appendix A Proof of Lemma 5
The following Lemma will be frequently used throughout the appendix.
@@
This completes the proof of (17) and the proof of (18) is similar.
@@
Proof  For any state $s$, we have
$\left|T^{\pi^{t+1}}_{I}\bar{V}^{t}(s)-T_{I}\bar{V}^{t}(s)\right|$
$=\left|\sum_{a}\pi^{t+1}(a|s)\left(\mathcal{R}(s,a)+\gamma\sum_{s^{\prime}}\bar{P}(s^{\prime}|s,a)\bar{V}^{t}(s^{\prime})\right)-T_{I}\bar{V}^{\pi^{t}}\right|$
=$\stackrel{{\scriptstyle(a)}}{{=}}\left|\sum_{n=1}^{N}q_{n}\left[\sum_{a}\pi^{t+1}_{n}(a|s)\left(\mathcal{R}(s,a)+\gamma\sum_{s^{\prime}}\bar{P}(s^{\prime}|s,a)\bar{V}^{t}(s^{\prime})\right)-T_{I}\bar{V}^{\pi^{t}}\right]\right|$
=$\stackrel{{\scriptstyle(b)}}{{=}}\left|\sum_{n=1}^{N}q_{n}\left[\gamma\sum_{a,s^{\prime}}\pi^{t+1}_{n}(a|s)\left(\bar{P}(s^{\prime}|s,a)-P_{n}(s^{\prime}|s,a)\right)\bar{V}^{t}(s^{\prime})+T^{\pi^{t+1}_{n}}_{n}\bar{V}^{t}(s)-T_{I}\bar{V}^{t}\right]\right|$
$\stackrel{{\scriptstyle(c)}}{{\leq}}\left|\sum_{n=1}^{N}q_{n}\gamma\sum_{a,s^{\prime}}\pi^{t+1}_{n}(a|s)\left(\bar{P}(s^{\prime}|s,a)-P_{n}(s^{\prime}|s,a)\right)\bar{V}^{t}(s^{\prime})\right|$
$\quad+\left|\sum_{n=1}^{N}q_{n}\left(T^{\pi^{t+1}_{n}}_{n}\bar{V}^{t}(s)-T_{n}\bar{V}^{t}(s)\right)\right|+\left|\sum_{n=1}^{N}q_{n}\left(T_{n}\bar{V}^{t}(s)-T_{I}\bar{V}^{t}(s)\right)\right|.$ (19)
Step (a) follows from $\pi^{t+1}(a|s)=\sum_{n=1}^{N}q_{n}\pi^{t+1}_{n}(a|s),\forall s\in\mathcal{S},a\in\mathcal{A}$. In step (b), we added and then subtracted the term $\sum_{a,s^{\prime}}\pi^{t+1}_{n}(a|s)P_{n}(s^{\prime}|s,a)\bar{V}^{t}(s^{\prime})$. The added term is combined with $\sum_{a,s^{\prime}}\pi^{t+1}_{n}(a|s)\mathcal{R}(s,a)$ to form $T^{\pi^{t+1}_{n}}_{n}\bar{V}^{t}(s)$. Step (c) follows from the triangle inequality.
@@
$\left|\sum_{n=1}^{N}q_{n}\gamma\sum_{a,s^{\prime}}\pi^{t+1}_{n}(a|s)\left(\bar{P}(s^{\prime}|s,a)-P_{n}(s^{\prime}|s,a)\right)\bar{V}^{t}(s^{\prime})\right|$
$=\left|\sum_{n=1}^{N}q_{n}\left(T^{\pi^{t+1}_{n}}_{I}\bar{V}^{t}(s)-T^{\pi^{t+1}_{n}}_{n}\bar{V}^{t}(s)\right)\right|$
@@
$\left|\sum_{n=1}^{N}q_{n}\left(T_{n}\bar{V}^{t}(s)-T_{I}\bar{V}^{t}(s)\right)\right|$ $\leq\frac{\gamma R_{\max}\kappa_{1}}{1-\gamma}.$ (21)
@@
$\left|T^{\pi^{t+1}}_{I}\bar{V}^{t}(s)-T_{I}\bar{V}^{t}(s)\right|$ $\leq\frac{2\gamma R_{\max}\kappa_{1}}{1-\gamma}+\bar{\epsilon}.$

### v6

@@ 3.5 Connection to Centralized Learning
When the environments are homogeneous, the policy is log-linear w.r.t. parameters, and the value function is linear w.r.t. parameters, the learning process of FAPI will be equivalent to learning from N copies of the same environment (that is identical to the imaginary MDP) with federated policy evaluation. Under such circumstances, Theorem 10 degenerates to that for centralized learning (6) (Bertsekas, 2022) with the same error bound, i.e., $\frac{\bar{\epsilon}-2\gamma\bar{\delta}}{(1-\gamma)^{2}}$, as shown in the following proposition.
@@
$\limsup_{t\to\infty}\left\|\bar{V}^{\pi^{t}}-\bar{V}^{\pi^{\ast}}\right\|$ $\leq\frac{\hat{\varepsilon}_{\theta}\sqrt{\left|\mathcal{A}\right|}R_{\max}}{\left(1-\gamma\right)^{3}}+\frac{2\gamma\bar{\varepsilon}_{w}}{\left(1-\gamma\right)^{2}}+\frac{\bar{\epsilon}+2\gamma\bar{\delta}}{(1-\gamma)^{2}},$
where $\hat{\varepsilon}_{\theta}$ is equal to $\bar{\varepsilon}_{\theta}$ and $\tilde{\varepsilon}_{\theta}$ for full participation and partial participation, respectively.
See Appendix F for the detailed proof. As discussed in Section 3.4, $\hat{\varepsilon}_{\theta}=0$ when the policy is log-linear w.r.t. parameters, and $\bar{\varepsilon}_{w}=0$ when the value function is linear w.r.t. parameters.
@@ 4.2 Implementation
To compute $\Delta_{n}$, we adopt the tabular maximum likelihood model (Moerland et al., 2020; Strehl et al., 2009; Ornik and Topcu, 2021; Strehl and Littman, 2008) in the implementation of FedPOHCS. Given a set of trajectories, let $C_{n}(s,a)$ denote the number of times action $a$ was taken under state $s$, $C_{n}(s,a,s^{\prime})$ represent the number of times the MDP transited from state $s$ to $s\prime$ after taking action $a$, and $r_{n}$ denote the corresponding sequence of reward received. Generally speaking, the more trajectories we have, the more accurate the modeling would be. For Mountain Cars and Hoppers, the number of trajectories is 200. For HongKongOSMs, the number of trajectories is 1000. We can estimate the transition probability $P_{n}$ of the $n$-th MDP and the reward function $\mathcal{R}$ by
@@
where the time horizon $t$ is a hyperparameter. Then, we diagonalize the vector $D_{n,s^{\prime}}$ to obtain the state visitation frequency matrix $\hat{\mathbf{D}}_{\pi,\mu,P_{n}}$. Last, we estimate the advantage function by Generalized Advantage Estimation (GAE) (Schulman et al., 2016).
@@ 5 Experiments
In this section, we introduce two federated environments for empirical evaluation and evaluate the proposed client selection method from different perspectives. All experimental results are reproducible and can be accessed from the GitHub repository: https://github.com/ShiehShieh/FedPOHCS.
@@ 5.1 Environments
We evaluate the effectiveness of the proposed client selection algorithm on a federated version of the mountain car continuous control problem (Moore, 1990) and the Mujoco Hopper problem (Coulom, 2002). Furthermore, we utilize the Flow simulator Vinitsky et al. (2018) and OpenStreetMap (OSM) (OpenStreetMap contributors, 2017; Vargas-Munoz et al., 2021) dataset to create a series of traffic networks for autonomous vehicle training. We construct the federated environments as follows:
@@
Hoppers consists of 60 equally weighted Hopper-v3 environments developed by OpenAI Gym. The environment has an 11-dimensional continuous state space and a 3-dimensional continuous action space. We introduce the heterogeneity into this system by following Jin et al. (2022), i.e., the leg size is uniformly sampled from $[0.01,0.07]$, $[0.01,0.10]$, and $[0.01,0.15]$ for low-level, medium-level, and high-level heterogeneity, respectively.
HongKongOSMs consists of 10 equally weighted traffic networks, each based on one of the OSM data as shown in Figure 1. Each traffic network contains one RL-controlled and 10 IDM-controlled (Intelligent Driver Model) vehicles. The 18-dimensional observation includes headway, speed, and positional information of visible neighborhoods of the RL-controlled vehicle, and only observations for vehicles running on the same and adjacent lanes are visible to the RL-controlled vehicle. The 2-dimensional action includes acceleration and lane-changing decisions. We adopted the Eclipse Simulation of Urban MObility (SUMO) simulator to conduct this experiment.
Figure 1: OSM data of ten areas in Hong Kong. (a) Causeway; (b) Central; (c) Chai Wan; (d) Clear Water Bay; (e) Kennedy Town; (f) Kai Tak; (g) North Point; (h) Po Lam; (i) Sham Shui Po; (j) Tseung Kwan O.
@@ 5.2 Experiment Settings
We use neural networks to represent policies as in Schulman et al. (2015); Vinitsky et al. (2018). Specifically, we use Multilayer Perceptrons (MLPs) with $\tanh$ non-linearity and hidden layers (64, 64). We use the SGD optimizer with a momentum of 0.9 and learning-rate decay of 0.98, 0.9, and 0.98 per round for Mountain Cars, Hoppers, and HongKongOSMs, respectively. Hyperparameters are carefully tuned so that they are near-optimal for each algorithm. See Appendix H for more experimental details.
We compare FedPOHCS with biased and unbiased client selection methods, including FedAvg (random selection), Power-of-Choice, and GradientNorm (Chen et al., 2021; Marnissi et al., 2021; Chen et al., 2022). FedAvg randomly selects K clients. Power-of-Choice utilizes the aforementioned two-phase scheme and selects candidate clients with the highest loss (or lowest advantages/values in case of RL problem). GradientNorm selects candidate clients with the largest gradient norm. Besides the client selection scheme, all algorithms follow the same procedure described at the beginning of Section 2.1. In particular, clients perform local training with the algorithm proposed by Schulman et al. (2017), with adaptive KL penalty term. At the end of every round, the server broadcasts the global policy to all clients, orders them to interact with their MDPs for several episodes (10 for Mountain Cars, 100 for Hoppers, and 20 for HongKongOSMs) and report the mean returns to evaluate the performance. Each experiment is averaged across three independent runs with different random seeds and parameter initializations, both of which are shared among all algorithms for a fair comparison. Confidence intervals are also reported.
Figure 2: Comparison of FedAvg, Power-of-Choice, FedPOHCS, and GradientNorm on Mountain Cars with a medium level of heterogeneity. For FedAvg, the learning rate is 0.005 and the KL target is 0.003. For Power-of-Choice, GradientNorm, and FedPOHCS, the learning rate is 1e-3, and the KL target is 0.003.
Figure 3: Comparison of FedAvg, Power-of-Choice, FedPOHCS, and GradientNorm on Hoppers with a medium level of heterogeneity. For all algorithms, the learning rate is 0.03, the learning rate decay is 0.9, and the KL target is 0.003.
Figure 4: Comparison of FedAvg, Power-of-Choice, FedPOHCS, and GradientNorm on HongKongOSMs. For all algorithms, the learning rate is 0.0001, the learning rate decay is 0.98, and the KL target is 0.0001.
@@ 5.3 Performance and Stability
Although the original mountain car problem is simple and most modern RL algorithms can easily obtain a score over 90.0, the federated setting imposes great difficulties in solving it. Figure 2 shows the performance comparison between FedPOHCS and several baselines on Mountain Cars with medium-level heterogeneity. It can be observed that FedPOHCS has a faster convergence speed and a more stable learning process. To compute the selection metric, we discretize the state and action by rounding them off to the nearest 0.1, resulting in about 8000 states and 100 actions that are frequently visited. In each round, the first phase of FedPOHCS takes 1-10 seconds, and the local training takes 20-30 seconds. Although the running time of the first phase may vary depending on the implementation, FedPOHCS outperforms all baselines in terms of the number of rounds and wall-clock time in our setting. In Figure 2, we have also included the communication-efficient variant, i.e., the one-phase scheme, for comparison purposes. It can be observed that the one-phase scheme may slow down and destabilize learning in the early stage of training, but the final performance is comparable to the two-phase scheme.
We can draw a similar conclusion on Hoppers with medium-level heterogeneity. As shown in Figure 3, FedPOHCS can obtain an accumulated reward of 1450 within 20 rounds of training, while it takes about 80 rounds for Power-of-Choice to reach 1450, demonstrating the advantage of the proposed FedPOHCS algorithm in speeding up convergence.
HongKongOSMs is much more difficult to solve as shown by the high-variance curves in Figure 4. Since different OSM datasets have different numbers of lanes, target velocity, and maximum acceleration/deceleration, their state spaces and transition probabilities may be highly distinct from each other. Compared with other approaches, FedPOHCS can obtain higher rewards and converge to the highest point.
@@ 6 Conclusion
This appendix contains the proof of Lemma 5 in Appendix A, the proof of Lemma 6 in Appendix B, the proof of Proposition 7 in Appendix C, the proof of Theorem 10 in Appendix D, the proof of Proposition 9 in Appendix E, the proof of Proposition 14 in Appendix F, and additional experiment setting in Appendix H.
@@ Appendix A Proof of Lemma 5
The following Lemmas will be frequently used throughout the appendix.
@@
This completes the proof of (24), and the proof of (25) is similar.
Lemma 16
Let $\bar{\pi}^{t}(a|s)=\sum_{n=1}^{N}q_{n}\pi^{t}_{n}(a|s)$ denote the expected output of all local policies and $\tilde{\pi}^{t}(a|s)=\sum_{m\in\mathcal{C}}q^{\prime}_{m}\pi^{t}_{m}(a|s),\forall s\in\mathcal{S},a\in\mathcal{A}$ represent the expected output of a set $\mathcal{C}$ of local policies. Define $\bar{\varepsilon}_{\theta}=\max_{t}\left\|\pi^{t}(\cdot|s)-\bar{\pi}^{t}_{n}(\cdot|s)\right\|_{2}$ and $\tilde{\varepsilon}_{\theta}=\max_{t}\left\|\pi^{t}(\cdot|s)-\tilde{\pi}^{t}_{m}(\cdot|s)\right\|_{2}$ for full participation and partial participation, respectively. For any value function $V\in\mathbb{R}^{\left|\mathcal{S}\right|}$ and policy $\pi^{t}$ at round $t$, we have
$\left\|T_{I}^{\pi^{t}}V-T^{\bar{\pi}^{t}}_{I}V\right\|$ $\leq\frac{\bar{\varepsilon}_{\theta}\sqrt{\left|\mathcal{A}\right|}R_{\max}}{1-\gamma},$ (26)
$\left\|T_{n}^{\pi^{t}}V-T^{\bar{\pi}^{t}}_{n}V\right\|$ $\leq\frac{\bar{\varepsilon}_{\theta}\sqrt{\left|\mathcal{A}\right|}R_{\max}}{1-\gamma},$ (27)
$\left\|T_{I}^{\pi^{t}}V-T^{\tilde{\pi}^{t}}_{I}V\right\|$ $\leq\frac{\tilde{\varepsilon}_{\theta}\sqrt{\left|\mathcal{A}\right|}R_{\max}}{1-\gamma},$ (28)
$\left\|T_{n}^{\pi^{t}}V-T^{\tilde{\pi}^{t}}_{n}V\right\|$ $\leq\frac{\tilde{\varepsilon}_{\theta}\sqrt{\left|\mathcal{A}\right|}R_{\max}}{1-\gamma}.$ (29)
Proof  For any state $s$, we have
$\left|T_{I}^{\pi^{t}}V^{t}(s)-T^{\bar{\pi}^{t}}_{I}V^{t}(s)\right|$ $=\left|\sum_{a}\sum_{n=1}^{N}q_{n}\left(\pi^{t}(a|s)-\pi^{t}_{n}(a|s)\right)\left(\mathcal{R}(s,a)+\gamma\sum_{s^{\prime}}\bar{P}(s^{\prime}|s,a)V^{t}(s^{\prime})\right)\right|$
$\leq\left\|\pi^{t}(\cdot|s)-\sum_{n=1}^{N}q_{n}\pi^{t}_{n}(\cdot|s)\right\|_{2}\left\|\mathcal{R}(s,\cdot)+\gamma\sum_{s^{\prime}}\bar{P}(s^{\prime}|s,\cdot)V^{t}(s^{\prime})\right\|_{2}$
$\leq\frac{\sqrt{\left|\mathcal{A}\right|}R_{\max}}{1-\gamma}\left\|\pi^{t}(\cdot|s)-\sum_{n=1}^{N}q_{n}\pi^{t}_{n}(\cdot|s)\right\|_{2}$
$\leq\frac{\bar{\varepsilon}_{\theta}\sqrt{\left|\mathcal{A}\right|}R_{\max}}{1-\gamma},$
which completes the proof of (26) and the proofs of (27), (28), and (29) are similar.
@@
Proof  Let $\bar{\pi}^{t}(a|s)=\sum_{n=1}^{N}q_{n}\pi^{t}_{n}(a|s)$ denote the expected output of all local policies. For any state $s$, we have
$\left|T_{I}^{\pi^{t+1}}V^{t}(s)-T_{I}V^{t}(s)\right|$ $\leq\left|T_{I}^{\pi^{t+1}}V^{t}(s)-T^{\bar{\pi}^{t+1}}_{I}V^{t}(s)\right|+\left|T^{\bar{\pi}^{t+1}}_{I}V^{t}(s)-T_{I}V^{t}(s)\right|.$ (30)
@@
$\left|T_{I}^{\pi^{t+1}}V^{t}(s)-T^{\bar{\pi}^{t+1}}_{I}V^{t}(s)\right|$ $\leq\frac{\bar{\varepsilon}_{\theta}\sqrt{\left|\mathcal{A}\right|}R_{\max}}{1-\gamma}.$ (31)
Now we bound the second term on the RHS of (30). For any state $s$, we have
$\left|T^{\bar{\pi}^{t+1}}_{I}V^{t}(s)-T_{I}V^{t}(s)\right|$
$=\left|\sum_{a}\bar{\pi}^{t+1}(a|s)\left(\mathcal{R}(s,a)+\gamma\sum_{s^{\prime}}\bar{P}(s^{\prime}|s,a)V^{t}(s^{\prime})\right)-T_{I}V^{t}\right|$
=$\stackrel{{\scriptstyle(a)}}{{=}}\left|\sum_{n=1}^{N}q_{n}\left[\sum_{a}\pi^{t+1}_{n}(a|s)\left(\mathcal{R}(s,a)+\gamma\sum_{s^{\prime}}\bar{P}(s^{\prime}|s,a)V^{t}(s^{\prime})\right)-T_{I}V^{t}\right]\right|$
=$\stackrel{{\scriptstyle(b)}}{{=}}\left|\sum_{n=1}^{N}q_{n}\left[\gamma\sum_{a,s^{\prime}}\pi^{t+1}_{n}(a|s)\left(\bar{P}(s^{\prime}|s,a)-P_{n}(s^{\prime}|s,a)\right)V^{t}(s^{\prime})+T^{\pi^{t+1}_{n}}_{n}V^{t}(s)-T_{I}V^{t}\right]\right|$
$\stackrel{{\scriptstyle(c)}}{{\leq}}\left|\sum_{n=1}^{N}q_{n}\gamma\sum_{a,s^{\prime}}\pi^{t+1}_{n}(a|s)\left(\bar{P}(s^{\prime}|s,a)-P_{n}(s^{\prime}|s,a)\right)V^{t}(s^{\prime})\right|$
$\quad+\left|\sum_{n=1}^{N}q_{n}\left(T^{\pi^{t+1}_{n}}_{n}V^{t}(s)-T_{n}V^{t}(s)\right)\right|+\left|\sum_{n=1}^{N}q_{n}\left(T_{n}V^{t}(s)-T_{I}V^{t}(s)\right)\right|.$ (32)
Step (a) follows from $\bar{\pi}^{t+1}(a|s)=\sum_{n=1}^{N}q_{n}\pi^{t+1}_{n}(a|s),\forall s\in\mathcal{S},a\in\mathcal{A}$. In step (b), we added and then subtracted the term $\sum_{a,s^{\prime}}\pi^{t+1}_{n}(a|s)P_{n}(s^{\prime}|s,a)V^{t}(s^{\prime})$. The added term is combined with $\sum_{a,s^{\prime}}\pi^{t+1}_{n}(a|s)\mathcal{R}(s,a)$ to form $T^{\pi^{t+1}_{n}}_{n}V^{t}(s)$. Step (c) follows from the triangle inequality.
By Lemma 15, the first term on the RHS of (32) is upper bounded by
$\left|\sum_{n=1}^{N}q_{n}\gamma\sum_{a,s^{\prime}}\pi^{t+1}_{n}(a|s)\left(\bar{P}(s^{\prime}|s,a)-P_{n}(s^{\prime}|s,a)\right)V^{t}(s^{\prime})\right|$
$=\left|\sum_{n=1}^{N}q_{n}\left(T^{\pi^{t+1}_{n}}_{I}V^{t}(s)-T^{\pi^{t+1}_{n}}_{n}V^{t}(s)\right)\right|$
@@
$\left|\sum_{n=1}^{N}q_{n}\left(T_{n}V^{t}(s)-T_{I}V^{t}(s)\right)\right|$ $\leq\frac{\gamma R_{\max}\kappa_{1}}{1-\gamma}.$ (34)
@@
$\left|T^{\pi^{t+1}}_{I}V^{t}(s)-T_{I}V^{t}(s)\right|$ $\leq\frac{2\gamma R_{\max}\kappa_{1}}{1-\gamma}+\bar{\epsilon}.$ (35)
By substituting (35) and (31) into (30), we can obtain
$\left|T_{I}^{\pi^{t+1}}V^{t}(s)-T_{I}V^{t}(s)\right|\leq\frac{\bar{\varepsilon}_{\theta}\sqrt{\left|\mathcal{A}\right|}R_{\max}}{1-\gamma}+\frac{2\gamma R_{\max}\kappa_{1}}{1-\gamma}+\bar{\epsilon}.$ (36)